OpenStudy (anonymous):

Which of the following is the solution to 17 - 2x = -11? a.-14 b.-3 c.3 d.14

OpenStudy (anonymous):

@AbdullahM @puniUni16

OpenStudy (abdullahm):

add 11 to both sides so 17 - 2x = -11 + 11 +11 --------------- 17 + 11 - 2x = 0 Simplify that

OpenStudy (abdullahm):

by simplify I mean add the 17+11 and tell me what you get xD

OpenStudy (anonymous):

replace the variable (the letter) into parenthesis in your case x so pretty much just do what I do below: 17-2(-14)=45 and its not the correct answer 17-2(-3)=23 nope! 17-2(3)=11 close but no... 17-2(14)=-11 yayyy D is your answer